OpenStudy (anonymous):

Please Help Express the complex number in trigonometric form. -2i

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (sidsiddhartha):

-2i=2(cos(3Pi/2)+i sin(3Pi/2))=2(cos 270° + i sin 270°)

OpenStudy (anonymous):

thanks uchiha

OpenStudy (anonymous):

:)

OpenStudy (sidsiddhartha):

lol no pro :)